# Coherence-Assisted Entanglement Activation During Open Evolution

### v2

@@ Abstract
The fact that in the course of an open evolution there is, in a vast majority of cases, a creation of system-environment entanglement out of the initial coherences, has been known for many years. However, how much entanglement can be created under general assumptions has not been analyzed yet. In this work, we investigate system-environment entanglement generation under a broad and important class of interactions known as pure dephasing interactions. Using completely general initials states, we derive a series of bounds on the amount of the created entanglement, connecting it to the initial system coherences as well as to the decohrence strength, making quantitatively precise the statement that larger entanglement leads to more decoherence. Interestingly, we obtain lower bounds, certifying minimal amount of entanglement that can be produced. As an illustration, we apply our results to the spin-boson model, where a qubit interacts with a bosonic bath, showing time evolution of the entanglement and the Markovian/non-Markovian crossover. Our results help to better understand and quantify the complicated correlations, produced during an open evolution, which will hopefully lead to better/new decoherence mitigation techniques.
1 Introduction
@@
Entanglement [1] has been the subject of quite amount of researches in last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well. Understandably, the vast majority of entanglement studies focus on the purposeful entanglement generation, manipulation, and preservation. However, there are processes where entanglement can be generated or activated spontaneously. Open systems dynamics (see e.g. [2, 3] for a modern introduction) is one of such examples, where the interaction between the system and its environment leads in general to entanglement creation out of initial system coherences.
@@
Until recently [4, 5, 6, 7] this system-environment entanglement has not been systematically studied using modern methods of quantum information theory, being mostly treated as a mere source of decoherence. However, with the emergence and the gradual acceptance of a new paradigm in open systems, given by the quantum Darwinism idea [8] and its refinements such as Spectrum Broadcast Structures and strong Darwinism (see e.g. [9] for a review), the environment has been recognized as a carrier of valuable information about the system. Therefore, studying system-environment entanglement is important not only for understanding decoherence processes, since entanglement in general increases decoherence efficiency, see e.g. [10] for recent take on this question, but also to know how environment acquires system-related information.
@@
One can illustrate the system-environment entanglement activation on a paradigmatic example of the, so called, pure dephasing evolution, also known as controlled-unitary evolution, alternatively called also a measurement-type evolution:
@@
while the reduced subsystem $\textrm{Tr}_{E}[|\psi^{\prime}_{SE}\rangle\langle\psi^{\prime}_{SE}|]$ lost the initial coherence in $|\psi_{S}\rangle$. In the modern language of the resource theories [14], this an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  [15, 16, 17, 18, 19, 20]. Let us remark here that an abstract coherence theory finds its natural application in open systems with pure dephasing evolution as the latter contains a natural, preferred basis.
@@
However easy the pure state interconversion (2) looks like, the situation gets much more complicated when considering mixed states, especially if one wants to quantify the amount of the generated entanglement. Assuming that the initial state of the system is pure, Roszak et al.  [4, 5] recently found necessary and sufficient separability conditions for the system-environment split and have developed own operational measure of the system-environment entanglement [6, 7]. However, it is still unknown how to generalize their methods to mixed initial states of the system.
@@
Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system respect to relative entropy measure [17]. However, for given Hamiltonian, the corresponding pure dephasing evolution is not incoherent operation generally and we also take the environmental state as arbitrary, their finding does not hold in our case. In pure dephasing evolution, therefore, it is not clear how entanglement generation is quantified through the initial coherence or decoherence in the system.
@@
In this work, we study entanglement generation for mixed states of the system. We derive two lower bounds and one upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ [21, 22]. Following the simple intuition (2) and its developments [15, 16, 17, 18, 19, 20], the first lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
@@
The second lower bound and the upper bound are quantified with the decoherence in the system, following the intuition that entanglement generation requires not only the initial coherence but the decoherence the system after the evolution which is also suggested by (2).
@@
Regarding the first lower bound, for two-dimensional central systems, i.e. for qubits, we derive a stronger result in given by a closed formula that links entanglement with the distinguishability of the environment states. In particular, this allows to study entanglement dynamics for open qubits, understood as the real-time monitoring of the minimum amount of entanglement produced during the evolution. As an illustration, we apply our results to the spin-boson model (see e.g. [2] and the references therein). We study the dependence of the minimum generated entanglement on various model parameters e.g. the temperature and Ohmicity. Varying the latter through the well-known Markovian/non-Markovian transition, we see the traces of non-Markovianity in the dynamics of entanglement; cf. [23].
2 Preliminaries
Pure dephasing evolution is defined by the following Hamiltonian:
$H=H_{S}+H_{E}+\sum_{i=0}^{N-1}|i\rangle\langle i|\otimes V_{i}.$ (3)
$H_{S}$ is the self Hamiltonian of the system which here we assume to be commuting with the interaction, $H_{S}=\sum_{i}\varepsilon_{i}|i\rangle\langle i|$. The case of a more general $H_{S}$ is out of the scope of this work as we concentrate on the pure dephasing models. $H_{E}$ and $V_{i}$’s are arbitrary environmental observables. (3) is the most general Hamiltonian that generates evolution of the type (2). The unitary evolution $U$ is given by:
$U(t)=\sum_{i}|i\rangle\langle i|\otimes U_{E}^{i}(t),\quad U_{E}^{i}(t)=e^{-(\varepsilon_{i}+H_{E}+V_{i})t},$ (4)
@@ 2 Preliminaries
which is the controlled-unitary or measurement-type operation in the quantum information terminology. The basis $|i\rangle$ is the pointer basis. Since observables $V_{i}$ are arbitrary, in general non-commuting, there does not exist a common diagonalizing basis, i.e. an incoherent basis for the environment, which makes the whole unitary operator incoherent. We will review the definition of incoherent operations below.
@@
We will assume that the system and the environment are initially product, i.e. $\rho_{SE}(0)=\rho_{S}\otimes\rho_{E}$. This assumption is justified here by the fact that we are interested in the build-up of correlations, thus we want to initiate the system in a completely uncorrelated state, so that any produced correlations are due to the interaction. Applying the unitary operator defined in (4), we obtain the following state $\sigma_{SE}$ after the interaction:
$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}$
$=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
@@
$E_{r}(\rho_{SE})=\min_{\xi_{SE}\in Sep}S(\rho_{SE}||\xi_{SE})$ (6)
where $S(\rho||\xi)=\textrm{Tr}[\rho(\log\rho-\log\xi)]$ is the quantum relative entropy and the minimization is taken over all separable states $\xi_{SE}=\sum_{i}p_{i}\xi_{S}^{i}\otimes\xi_{E}^{i}$. This is a proper measure of entanglement [24], with a clear operational interpretation as the exponential bound on asymptotic probability of error of confusing $\rho_{AB}$ with the nearest separable state via POVM measurements, see e.g. [25]. The minimization in definition is the reason why it is difficult to evaluate the value directly. We employ the following lemma to bypass this problem:
@@
where $S(\rho)$ is the von Neumann entropy: $S(\rho)=-\textrm{Tr}[\rho\log\rho]$.
We will also need elements of the quantum coherence theory; see e.g.  [27]. Since quantum coherence is a basis dependent resource, we need to specify which basis is incoherent. In principle, one can develop a theory with any incoherent basis. However, since we study a concrete family of open dynamics, given by (3), the pointer basis of the central system $|i\rangle$ is the natural candidate here because if the system loses coherence due to the interaction, decoheres to this basis. We do not specify the incoherent basis for environment because the coherence in the environment does not matter in our analysis as we will show below. In the language of the coherence theory, an incoherent state is simply a diagonal state in the incoherent basis, which is of course nothing else but a decohered state in the open systems/quantum foundations terminology. We denote by $\mathcal{I}$ the set of all incoherent states. In a resource theory, of which coherence theory is an example, an operation which does not create resource is called a free operation respect to that resource. Various free operations for coherence are proposed for different physical motivations [28]. The most standard one is incoherent operation whose Kraus operators $\{K_{n}\}$ satisfy $\frac{K_{n}\rho K_{n}^{\dagger}}{\textrm{Tr}[K_{n}\rho K_{n}^{\dagger}]}\in\mathcal{I}$ for all $n$ and any incoherent state $\rho\in\mathcal{I}$. One can now define the relative entropy of coherence $C_{r}$ in analogy to the relative entropy of entanglement (6) [27]:
@@
$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{i}\}}\sum_{i}\sqrt{\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]}$
@@
where $\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]$ is the probability that one obtains the same value $i$ from the measurement on $\rho$ and $\sigma$. Vanishing state fidelity is equivalent to a perfect distinguishability of the two states by POVM’s. This justifies the use of the state fidelity as a measure of distinguishability. This measure does not have as clear operational meaning as e.g. the quantum Chernoff bound [30], related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (2) and it also provides both an upper and a lower bound for the quantum Chernoff information.
Discriminating states from an ensemble =$\mathcal{E}=\{p_{i},\rho_{i}\}$ is a more difficult task and its efficiency can be quantified by the, so called, accessible information $I_{acc}(\mathcal{E})$. It is defined as the maximum of the Shannon (classical) mutual information $I(i:m)$ [31] between the ensemble index $i$ and the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$. The formal expression is given by:
$I_{acc}=\max_{POVM:\{M_{m}\}}I(i:m),$ (10)
where $I(i:m)$ is defined using the conditional probabilities of measuring $m$ while the sate was $\rho_{i}$: $p(m|i)=\textrm{Tr}(M_{m}\rho_{i})$.
@@
[32][Lemma 3.6] Let =$\mathcal{E}=\{p,\rho_{0},(1-p),\rho_{1}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
@@
where $h$ is the binary Shannon entropy: $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$.
3 Bounds on the system-environment entanglement generation
@@
$\sigma_{E}(t)=\textrm{Tr}_{S}[U(t)\rho_{S}\otimes\rho_{E}U(t)^{\dagger}]=\sum_{i}p_{i}\rho_{i}(t),$ (12)
where $p_{i}=\langle i|\rho_{S}|i\rangle$ are the initial pointer probabilities and the conditional states of the environment $\rho_{i}(t)=U_{E}^{i}(t)\rho_{E}{U_{E}^{i}}(t)^{\dagger}$ encode, in a generally imperfect way, the information about the state of the central system. One can now try to recover this information by implementing a POVM measurement $\{M_{m}\}$ on the environment in order to extract the value of $i$, encoded by the environment in the course of the evolution. One may view this as an indirect measurement of the system through the environment, common in the macroscopic world; cf. the studies of objectivity in quantum mechanics in e.g. [8, 9]. The uncertainty of this indirect measurement is represented by Shannnon conditional entropy $H(i|m)$ [31]. From above observations, it is natural to think the less this uncertainty is, the stronger the system and the environment are correlated. On the other hand, the more is there initial coherence in the pointer basis $\{|i\rangle\}$, the more entanglement should be generated [18]. Indeed this is so as the following Lemma shows:
@@
Applying Lemma 1 to $\sigma_{SE}=U\rho_{S}\otimes\rho_{E}U^{\dagger}$, we have
[…]
$S(\sigma_{E})-S(\sigma_{SE})\}.$ (14)
We focus on finding a lower bound for $S(\sigma_{E})-S(\sigma_{SE})$ because $S(\sigma_{E})$ can be transformed to the conditional entropy or the overlap of environmental states through lemma 2, leading us to a nice result represented in corollary 1. We will also consider a lower bound for $S(\sigma_{S})-S(\sigma_{SE})$ later. We first observe that:
@@
where we used fact that the unitary evolution preserves the von Neumann entropy and it is additive for product states. Notice (15) can be negative while the relative entanglement should be always non-negative.
Next, let us recall the Holevo bound on the $I_{acc}$ associated with (12):
$I_{acc}\leq S(\sigma_{E})-\sum_{i}\langle i|\rho_{S}|i\rangle S(U_{i}\rho_{E}U_{i}^{\dagger}),$ (16)
@@
since $S(U_{i}\rho_{E}U_{i}^{\dagger})=S(\rho_{E})$. From the definition of the Shannon mutual information $I(x:y)=H(x)-H(x|y)$, where $H(x)$ and $H(x|y)$ are the classical Shannon and relative entropies respectively, the accessible information $I_{acc}$ can be represented as follows.
$I_{acc}=$ $\max_{POVM:\{M_{m}\}}I(i:m)$
$=$ $\max_{POVM:\{M_{m}\}}\left[H(i)-H(i|m)\right].$ (18)
Index $i$ is a random variable, distributed with the initial pointer probabilities $p_{i}$, as follows from (12), or equivalently with the eigenvalues of the fully decohered state $\Delta(\rho_{S})=\sum_{i}|i\rangle\langle i|\rho_{S}|i\rangle\langle i|$. Therefore we may write $H(i)=S(\Delta(\rho_{S}))$. Since there is a maximization over all POVM in the definition (3), for an arbitrary POVM $\{M_{m}\}$ we will have:
$I_{acc}\geq H(i)-H(i|m)=S(\Delta(\rho_{S}))-H(i|m)$ (19)
@@
$\geq$ $S(\Delta(\rho_{S}))-H(i|m)-S(\rho_{S})$
$=$ $C_{r}(\rho_{S})-H(i|m),$
@@
The above Lemma is valid for any finete dimensionality of the central system and the environment. The price to pay for this generality is the presence of the arbitrary POVM in the bound (15). This arbitrariness can be removed when the central system is a qubit. Then, there ensemble (12) consists of only two states $\rho_{0}(t)=U_{0}(t)\rho_{E}U_{0}(t)^{\dagger}$, $\rho_{1}(t)=U_{1}(t)\rho_{E}U_{1}(t)^{\dagger}$ with probabilities $p=\langle 0|\rho_{S}|0\rangle$ and $1-p=\langle 1|\rho_{S}|1\rangle$. We can then apply the Lemma 2 and obtain:

### v5

@@ Abstract
The fact that in the course of an open evolution there is, in a vast majority of cases, the creation of system-environment entanglement out of the initial coherence, has been known for many years. However, how much entanglement can be created under general assumptions has not been analyzed yet. In this work, we investigate system-environment entanglement generation under a broad and important class of interactions known as pure dephasing interactions and for arbitrary initial conditions. Our main results are the following bounds on the relative entropy of entanglement $E_{r}(\sigma_{SE})$ after evolution: $C_{r}(\rho_{S})-H(\mathcal{I}|\mathcal{M})\leq E_{r}(\sigma_{SE})\leq C_{r}(\rho_{S})-C_{r}(\sigma_{S})+H(\mathcal{I})$, connecting it on one side to the initial relative entropy of coherence, $C_{r}(\rho_{S})$, and information extractable from the environment, $H(\mathcal{I}|\mathcal{M})$, and on the other side to the decoherence strength, $C_{r}(\rho_{S})-C_{r}(\sigma_{S})$, making quantitatively precise the statement that larger entanglement leads to more decoherence. As an illustration, we apply our results to the spin-boson model, where a qubit interacts with a bosonic bath, showing time evolution of the entanglement and the Markovian/non-Markovian crossover. Our results help better understand and quantify the complicated correlations, produced during an open evolution, which will hopefully lead to better and new decoherence mitigation techniques.
I Introduction
@@
Entanglement [1] has been the subject of a significant amount of research in the last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well. Understandably, the vast majority of entanglement studies focus on the purposeful entanglement generation, manipulation, and preservation. However, there are processes where entanglement can be generated or activated spontaneously. Open system dynamics (see e.g. [2, 3] for a modern introduction) is one such example, where the interaction between the system and its environment leads in general to entanglement creation out of initial system coherences.
@@
Until recently [4, 5, 6, 7], this system-environment entanglement had not been systematically studied using modern methods of quantum information theory, having been mostly treated as a mere source of decoherence. However, with the emergence and the gradual acceptance of a new paradigm in open systems, given by the quantum Darwinism idea [8] and its refinements such as Spectrum Broadcast Structures and strong Darwinism (see e.g. [9] for a review), the environment has been recognized as a carrier of valuable information about the system. Therefore, studying system-environment entanglement is important not only for understanding decoherence processes, since entanglement in general increases decoherence efficiency, see e.g. [10] for recent take on this question, but also to know how environment acquires system-related information.
@@
One can illustrate the system-environment entanglement activation on a paradigmatic example of the so-called pure dephasing evolution, also known as controlled-unitary evolution, alternatively called also a measurement-type evolution:
@@
while the reduced subsystem $\textrm{Tr}_{E}[|\psi^{\prime}_{SE}\rangle\langle\psi^{\prime}_{SE}|]$ lost the initial coherence in $|\psi_{S}\rangle$. In the modern language of the resource theories [14], this is an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  [15, 16, 17, 18, 19, 20]. Let us remark here that an abstract coherence theory naturally applies to open systems with pure dephasing evolution as the latter contains a natural, preferred basis.
@@
Although the pure state interconversion (2) appears simple, the situation gets much more complicated when considering mixed states, especially if one wants to quantify the amount of the generated entanglement. Assuming that the initial state of the system is pure, Roszak et al.  [4, 5] recently found necessary and sufficient separability conditions for the system-environment split and have developed an operational measure of the system-environment entanglement [6, 7]. However, it is still unknown how to generalize their methods to mixed initial states of the system.
@@
We introduce some terminology for readers not familiar with quantum coherence. Roughly speaking, an incoherent state is simply a state that is diagonal with respect to some reference basis, known as the incoherent basis, and an incoherent operation is a quantum operation that maps incoherent states to incoherent states. Detailed definitions will be reviewed later. Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system with respect to relative entropy measure [17]. However, for given Hamiltonian, the corresponding pure dephasing evolution is not incoherent operation generally and we also take the environmental state as arbitrary, their finding does not hold in our case. In pure dephasing evolution, therefore, it is not clear how entanglement generation is quantified through the initial coherence or decoherence in the system.
@@
In this work, we study entanglement generation for mixed states of the system. We derive a lower bound and an upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ [21, 22]. Following the simple intuition (2) and its developments [15, 16, 17, 18, 19, 20], the lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
@@
The upper bound is quantified with the strength of decoherence in the system, following the intuition that entanglement generation requires not only the initial coherence but the interaction with the environment followed by decoherence which is also suggested by (2).
@@
Regarding the first lower bound, for two-dimensional central systems, i.e. for qubits, we derive a stronger result in given by a closed formula that links entanglement with the distinguishability of the environment states. In particular, this allows us to study entanglement dynamics for open qubits, understood as the real-time monitoring of the minimum amount of entanglement produced during the evolution. As an illustration, we apply our results to the spin-boson model (see e.g. [2] and the references therein). We study the dependence of the minimum generated entanglement on various model parameters e.g. the temperature and Ohmicity. Varying the latter through the well-known Markovian/non-Markovian transition, we see the traces of non-Markovianity in the dynamics of entanglement; cf. [23].
II Preliminaries
We consider here the following Hamiltonian:
$H_{tot}=H_{S}+H_{E}+\sum_{i=0}^{N-1}|i\rangle\langle i|\otimes V_{i}$ (3)
where $H_{S}$ is the self-Hamiltonian of the system which we assume to commute with the interaction, i.e. $H_{S}=\sum_{i}\varepsilon_{i}|i\rangle\langle i|$. The whole resulting Hamiltonian $H_{tot}$ then takes the so-called pure dephasing form: $H_{tot}=\sum_{i}|i\rangle\langle i|\otimes(\varepsilon_{i}+V_{i}+H_{E})\equiv\sum_{i}|i\rangle\langle i|\otimes V_{i}^{\prime}$, with $V_{i}^{\prime}\equiv\varepsilon_{i}+V_{i}+H_{E}$. This Hamiltonian preserves the pointer basis, i.e. $|i\rangle\langle i|,H_{tot}]=0$, thus affecting only the off-diagonal terms, i.e. the coherence, via the dephasing noise, hence the name. We make this choice for simplicity, since inclusion of a general $H_{S}$ complicates the dynamics immensely, which forces to use approximate methods such as master equations [2, 3] and ultimately the numerical methods. This type of a much more advanced and complicated analysis is out of the scope of the current work and is left for a future research. Here, $H_{E}$ and $V_{i}$’s are arbitrary environmental observables. Eq. (3) is the most general Hamiltonian that generates evolution of the type (2). The unitary evolution $U$ is given by:
$U(t)=\sum_{i}|i\rangle\langle i|\otimes U_{E}^{i}(t),\quad U_{E}^{i}(t)=e^{-j(\varepsilon_{i}+H_{E}+V_{i})t},$ (4)
@@ II Preliminaries
where $j$ is imaginary unit. This unitary is also called the controlled-unitary or measurement-type operation in the quantum information terminology. The basis $|i\rangle$ is the pointer basis. Since observables $V_{i}$ are arbitrary, in general non-commuting, there does not exist a common diagonalizing basis, i.e. an incoherent basis for the environment, which makes the whole unitary operator incoherent. We will review the definition of incoherent operations below.
@@
We will assume that the system and the environment are initially in a product state, i.e. $\rho_{SE}(0)=\rho_{S}\otimes\rho_{E}$. This assumption is justified here by the fact that we are interested in the build-up of correlations, thus we want to initiate the system in a completely uncorrelated state, so that any produced correlations are due to the interaction. Applying the unitary operator defined in (4), we obtain the following state $\sigma_{SE}$ after the interaction:
$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
@@
$E_{r}(\rho_{SE})=\min_{\xi_{SE}\in\textrm{Sep}}S(\rho_{SE}||\xi_{SE})$ (6)
where $S(\rho||\xi)=\textrm{Tr}[\rho(\log\rho-\log\xi)]$ is the quantum relative entropy and the minimization is taken over all separable states $\xi_{SE}=\sum_{i}p_{i}\xi_{S}^{i}\otimes\xi_{E}^{i}$. Throughout this paper the base used in the logarithm is 2. This is a proper measure of entanglement [24], with a clear operational interpretation as the exponential bound on asymptotic probability of error of confusing $\rho_{AB}$ with the nearest separable state via POVM measurements, see e.g. [25]. The minimization in definition is the reason why it is difficult to evaluate the value directly. We employ the following lemma to bypass this problem:
@@
where $S(\rho)$ is the von Neumann entropy: $S(\rho)=-\textrm{Tr}[\rho\log\rho]$[27].
We will also need elements of the quantum coherence theory; see e.g.  [28]. Since quantum coherence is a basis dependent resource, we need to specify which basis is incoherent. In principle, one can develop a theory with any incoherent basis. However, since we study a concrete family of open dynamics, given by (3), the pointer basis of the central system $|i\rangle$ is the natural candidate here because if the system loses coherence due to the interaction, decoheres to this basis. We do not specify the incoherent basis for environment because the coherence in the environment does not matter in our analysis as we will show below. In the language of coherence theory, an incoherent state is simply a diagonal state in the incoherent basis, which is of course nothing else but a decohered state in the open systems/quantum foundations terminology. In a resource theory, of which coherence theory is an example, an operation which does not create resource is called a free operation with respect to that resource. Various free operations for coherence are proposed for different physical motivations [29]. The most standard one is incoherent operation whose Kraus operators $\{K_{n}\}$ satisfy that $\frac{K_{n}\rho K_{n}^{\dagger}}{\textrm{Tr}[K_{n}\rho K_{n}^{\dagger}]}$ is incoherent for all $n$ and any incoherent state $\rho$. One can now define the relative entropy of coherence $C_{r}$ in analogy to the relative entropy of entanglement (6) [28]:
@@
$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{m}\}}\sum_{m}\sqrt{\textrm{Tr}[\rho M_{m}]\textrm{Tr}[\sigma M_{m}]}$
@@
where $\textrm{Tr}[\rho M_{m}]\textrm{Tr}[\sigma M_{m}]$ is the probability that one obtains the same value $m$ from the measurement on $\rho$ and $\sigma$. Vanishing state fidelity is equivalent to a perfect distinguishability of the two states by POVM’s. This justifies the use of the state fidelity as a measure of distinguishability. This measure does not have as clear operational meaning as e.g. the quantum Chernoff bound [31], related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (II) and it also provides both an upper and a lower bound for the quantum Chernoff information.
Discriminating states from an ensemble $\mathcal{E}=\{p_{i},\rho_{i}\}_{i}$ is a more difficult task and its efficiency can be quantified by the, so-called, accessible information $I_{acc}(\mathcal{E})$. It is defined as the maximum of the Shannon (classical) mutual information $I(\mathcal{I}:\mathcal{M})$ [32] between the variable $\mathcal{I}$ for the ensemble index $i$ and the variable $\mathcal{M}$ for the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$. Recall that mutual information is defined as I(X:Y)=H(X)H(X|Y)I(X:Y)=H(X)-H(X|Y) where $H(X)$ is the Shannon entropy and $H(X|Y)$ is the conditional entropy for random variable $X$ and $Y$ [32]. The formal expression of accessible information is given by:
$I_{acc}=$ $\max_{POVM:\{M_{m}\}}I(\mathcal{I}:\mathcal{M})=\max_{POVM:\{M_{m}\}}H(\mathcal{I})-H(\mathcal{I}|\mathcal{M}),$ (10)
where the conditional entropy $H(\mathcal{I}|\mathcal{M})$, associated with the POVM $\{M_{m}\}$ is defined as:
$H(\mathcal{I}|\mathcal{M})=-\sum_{i,m}p_{i}\textrm{Tr}[\rho_{i}M_{m}]\log\frac{p_{i}\textrm{Tr}[\rho_{i}M_{m}]}{\textrm{Tr}[\sum_{j}p_{j}\rho_{j}M_{m}]}.$ (11)
@@
[33][Lemma 3.6] Let $\mathcal{E}=\{\{(p,\rho_{0}\},\{(1-p),\rho_{1}\}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
@@
where $h$ is the binary Shannon entropy: $h(p)=-p\log p-(1-p)\log(1-p)$.
III Bounds on the system-environment entanglement generation
@@
$\sigma_{E}(t)=\textrm{Tr}_{S}[\sigma_{SE}(t)]=\textrm{Tr}_{S}[U(t)\rho_{S}\otimes\rho_{E}U(t)^{\dagger}]=\sum_{i}p_{i}\rho_{i}(t),$ (13)
where $p_{i}=\langle i|\rho_{S}|i\rangle$ are the initial pointer probabilities and the conditional states of the environment $\rho_{i}(t)=U_{E}^{i}(t)\rho_{E}{U_{E}^{i}}(t)^{\dagger}$ encode, in a generally imperfect way, the information about the state of the central system. One can now try to recover this information by implementing a POVM measurement $\{M_{m}\}$ on the environment in order to extract the value of $\mathcal{I}$, encoded by the environment in the course of the evolution. One may view this as an indirect measurement of the system through the environment, common in the macroscopic world; cf. the studies of objectivity in quantum mechanics in e.g. [8, 9]. The uncertainty of this indirect measurement is represented by Shannon conditional entropy $H(\mathcal{I}|\mathcal{M})$ [32] as $\mathcal{I}$ represents the variable for index on the right hand side of (13) and $\mathcal{M}$ represents the variable for measurement value. From above observations, it is natural to think the less this uncertainty is, the stronger the system and the environment are correlated. On the other hand, the more is there initial coherence in the pointer basis $\{|i\rangle\}$, the more entanglement should be generated [18]. Indeed this is so as the following Lemma shows:
@@
Applying Lemma 1 to $\sigma_{SE}$, we have
$E_{r}(\sigma_{SE})\geq\max\{S(\sigma_{S})-S(\sigma_{SE}),S(\sigma_{E})-S(\sigma_{SE})\},$ (15)
where $\sigma_{S}=\textrm{Tr}_{E}[\sigma_{SE}]$. We focus on finding a lower bound for the latter, $S(\sigma_{E})-S(\sigma_{SE})$, because $S(\sigma_{E})$ can be transformed to the conditional entropy or the overlap of environmental states through lemma 2, leading us to a nice result represented in corollary 1. We first observe that:
@@
where we used fact that the unitary evolution preserves the von Neumann entropy and it is additive for product states. Notice (16) can be negative while the relative entropy of entanglement should be always non-negative.
Next, to find a lower bound of $S(\sigma_{E})$, let us recall the Holevo bound [34] on the $I_{acc}$ associated with (13):
$I_{acc}(\sigma_{E})\leq S(\sigma_{E})-\sum_{i}\langle i|\rho_{S}|i\rangle S(U_{i}\rho_{E}U_{i}^{\dagger}),$ (17)
@@
since $S(U_{i}\rho_{E}U_{i}^{\dagger})=S(\rho_{E})$. From the definition (10), $I_{acc}(\sigma_{E})$ is also bounded by $H(\mathcal{I})-H(\mathcal{I}|\mathcal{M})$. $\mathcal{I}$ is distributed with the initial pointer probabilities $p_{i}$, as follows from (13), or equivalently with the eigenvalues of the fully decohered state of system $\Delta(\rho_{S})=\sum_{i}|i\rangle\langle i|\rho_{S}|i\rangle\langle i|$. Therefore we may write $H(\mathcal{I})=S(\Delta(\rho_{S}))$. Since there is a maximization over all POVM in the definition (10), for an arbitrary POVM $\{M_{m}\}$ we will have:
$I_{acc}(\sigma_{E})\geq H(\mathcal{I})-H(\mathcal{I}|\mathcal{M})=S(\Delta(\rho_{S}))-H(\mathcal{I}|\mathcal{M})$ (19)
@@
$\geq$ $S(\Delta(\rho_{S}))-H(\mathcal{I}|\mathcal{M})-S(\rho_{S})$
$=$ $C_{r}(\rho_{S})-H(\mathcal{I}|\mathcal{M}),$
@@
The above Lemma is valid for any finite dimensionality of the central system and the environment. The price to pay for this generality is the presence of the arbitrary POVM in the bound (14). This arbitrariness can be removed when the central system is a qubit. Then, their ensemble (13) consists of only two states $\rho_{0}(t)=U_{0}(t)\rho_{E}U_{0}(t)^{\dagger}$, $\rho_{1}(t)=U_{1}(t)\rho_{E}U_{1}(t)^{\dagger}$ with probabilities $p=\langle 0|\rho_{S}|0\rangle$ and $1-p=\langle 1|\rho_{S}|1\rangle$. We can then apply the Lemma 2 and obtain: